OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. A. y - 6= -3x - 3 B. y - 6= -3(x + 1) C. y= -3x + 3 D. 3x + y= 3

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)